### v7

@@ 4.1 Preliminary Lemmas
Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete deterministic maximal $L^{p}$-regularity estimate is valid:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
@@
Additionally, for the discrete deterministic convolution operator $S_{0,h}$, we establish the following property.
Lemma 4.9.
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following inequality holds for all $\epsilon>0$:
$\|S_{0,h}g_{h}(t)\|_{C([0,T];\dot{H}_{h}^{2-2/p-\epsilon,q})}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
Proof.
For any $t\in[0,T]$, from the identity $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we deduce that
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$ $\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
where step (i) uses the estimate Eq. 4.21 with $\alpha=2-2/p-\epsilon$, and step (ii) applies Hölder’s inequality. Since the integral
$\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
is uniformly bounded with respect to $t\in[0,T]$, the desired inequality follows, thereby completing the proof of the lemma. ∎
@@
Suppose that $v\in L^{\infty}$. Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.11), and the function $G$ defined by equation (2.10). Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}v$ for $t\in[0,T]$. For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. For any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
@@
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.28)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.29)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.30)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\varepsilon}.$ (4.31)
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.33)
Here, $\beta^{\prime}$ denotes the conjugate exponent of $\beta$.
@@
The inequality Eq. 4.30 is derived in a similarly manner:
@@
The inequality Eq. 4.31 is straightforward to verify by applying Hölder’s inequality, in conjunction with the approximation property of $P_{h}$ from Eq. 4.19 and the regularity assertions in Theorem 3.1.
Now let us prove Eq. 4.32. By invoking Eq. 4.18 and replicating the reasoning behind [19, Equation (2.2)], we establish that
@@
$\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
$\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
Invoking the inequality Eq. 4.27 with $p$ replaced by $4p\beta/(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ belongs to $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can replicate the previous argument to conclude that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
Hence, Eq. 4.32 follows by combining the aforementioned inequalities. To deduce Eq. 4.33, we rely on Eq. 4.29 and proceed analogously to the proof of Eq. 4.32. This completes the proof. ∎
Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
@@
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (4.36)
From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
@@
Similarly, from Eq. 2.11 and the fact $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
@@
Defining $e_{h}:=z_{h}-P_{h}z$, we then obtain
@@
Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0,$
we use Eq. 4.37 to conclude that
@@
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
@@
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
@@
$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
@@
Regarding $I_{2}$, by Hölder’s inequality, Young’s inequality, and the inequalities Eq. 4.27 and Eq. 4.29, we can deduce that
$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.40)
For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.18, a straightforward calculation yields
@@
$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s\bigg{]}.$
By utilizing Eq. 4.31, we arrive at the following bound:
@@
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.34) and (4.35), along with the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we deduce that $P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}$. For any $s\in[0,t]$, we can then write:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}\mu$
@@
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.18), we obtain the following bound for $I_{4}$:
@@
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
Using (4.32) and (4.33), we further simplifies this bound to:
@@
By selecting $\beta$ sufficiently small within $\big{(}1,\frac{3p}{3p-4}\big{)}$ for any $\epsilon>0$, we arrive at the final bound for $I_{4}$:
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s.$ (4.42)
Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.18, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
@@
Subsequently, by Young’s inequality, we obtain
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}.$
Recalling the definitions in Eqs. 4.34 and 4.35 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.43)
which, in conjunction with Eqs. 4.27, 4.28 and 4.29, yields
$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
Consequently, we obtain the following bound for $I_{5}+I_{6}$:
@@
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.39, 4.40, 4.41, 4.42 and 4.44, we obtain the following inequality:

### v8

@@ 4.1 Preliminary Lemmas
Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following discrete maximal $L^{p}$-regularity estimate is valid:
$\lVert{S_{0,h}g_{h}}\rVert_{C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1-1/p,p}}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
@@
Remark 4.3.
A proof of Lemma 4.10 can be found in [12, Theorem 3.2]. Since the boundedness constant of the $H^{\infty}$-calculus for $-\Delta_{h}$ is independent of the spatial mesh size $h$, as shown in [19, Theorem 3.1], Lemma 4.11 follows directly from [32, Theorem 3.5]; see [19, Theorem 3.2] for further details.
@@
Suppose that $y_{0}\in L^{\infty}$ is deterministic. Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.9), and the function $G$ defined by equation (2.8). Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0}$ for $t\in[0,T]$. For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. For any $p,q\in(2,\infty)$ and $\epsilon>0$, the following inequalities hold:
@@
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.34)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.35)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\epsilon},$ (4.36)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\epsilon}.$ (4.37)
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\epsilon}.$ (4.39)
Here, $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$.
@@
$=c\lVert{(P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h})\Delta z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{0,q})}$
@@
$=ch^{1+2/p-\epsilon}\lVert{z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q})}$
[…]
The inequality Eq. 4.36 is derived in a similar manner:
@@
[…]
For Eq. 4.37, we apply Hölder’s inequality with $\alpha\in(1,2)$, yielding
$\mathbb{E}\left[\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\right]$
[…]
Choosing $\alpha>1$ sufficiently close to 1 so that $\frac{2}{\alpha}-\frac{\epsilon}{2}\geqslant 2-\epsilon$, and using the regularity estimate (3.3), we obtain (4.37).
Now, let us prove Eq. 4.38. Fix any $\beta\in(1,\frac{3p}{3p-4})$. By using Eq. 4.19 and following the reasoning behind [21, Equation (2.2)], we establish that
@@
[…]
Using Eq. 4.33 with $p$ replaced by $\frac{4p\beta}{4\beta-(3\beta-3)p}$ and $q$ by $2\beta^{\prime}$, we conclude that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
Moreover, by Theorem 3.1, we have $y\in L^{\frac{4\beta}{3(\beta-1)}}(\Omega\times(0,T);L^{2\beta})$, so a similar argument yields
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
Combining both estimates gives Eq. 4.38. The estimate Eq. 4.39 follows by a similar argument using Eq. 4.35. This completes the proof.
Finally, we are in a position to prove Theorems 3.2 and 3.3.
Proof of Theorem 3.2. Fix any $p\in(2,\infty)$. Define
@@
$G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0},\quad t\in[0,T].$ (4.42)
From Eq. 1.1 and the fact that $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
@@
Similarly, from Eq. 2.9 and the fact that $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
@@
Define $e_{h}:=z_{h}-P_{h}z$. Then we obtain
@@
Now define $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and consider the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0.$
Using Eq. 4.43, we conclude that
@@
By applying Itô’s formula to $\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ and using the initial condition $\eta_{h}(0)=0$ together with the definition of $F$ given by Eq. 2.4, we deduce that for any $0\leqslant t\leqslant T$,
@@
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}x\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
@@
$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
Next, we estimate the terms $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we obtain
@@
Regarding $I_{2}$, applying Hölder’s inequality, Young’s inequality, and the bounds given in Eq. 4.33 and Eq. 4.35, we deduce that
$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.46)
For $I_{3}$, using Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ stated in Eq. 4.19, a straightforward computation gives
@@
$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\right)\,\mathrm{d}s\bigg{]}.$
Using the estimate Eq. 4.37, we then arrive at the bound
@@
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.40) and (4.41), together with the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we find that
$P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}.$
Thus, for any $s\in[0,t]$, we compute:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}x$
@@
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.19), we derive the following estimate for $I_{4}$:
@@
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right),$
where $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$. Using the estimates (4.38) and (4.39), this simplifies further to
@@
By choosing $\beta$ sufficiently close to 1 within the interval $\left(1,\frac{3p}{3p-4}\right)$, for any $\epsilon>0$, we finally obtain
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.48)
Now, we proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ given in Eq. 4.19, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Invoking the condition Eq. 2.3, we further deduce that
@@
Applying Young’s inequality, we then have
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\right)\,\mathrm{d}s\bigg{]}.$
Recalling the definitions in Eqs. 4.40 and 4.41 and using the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T].$ (4.49)
Combining this with the estimates Eq. 4.33, Eq. 4.34, and Eq. 4.35, we find that
$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
As a result, we arrive at the following bound for $I_{5}+I_{6}$:
@@
Finally, by combining the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as established in Eqs. 4.45, 4.46, 4.47, 4.48 and 4.50, we obtain